OpenStudy (anonymous):

Radon enters through the soil into home basements, where it presents a health hazard if inhaled. In the simplest case of radon detection, a sample of air with volume V is taken. After equilibrium has been established, the radioactive decay D of the radon gas is counted with efficiency E over time t. The radon concentration C present in the sample of air varies directly as the product of D and E and inversely as the product of V and t. For a fixed radon concentration C and time t, find the change in the radioactive decay count D if V is doubled and E is reduced by 50%.

OpenStudy (anonymous):

D increases by _____%

OpenStudy (richyw):

holy cow this one has a lot of letters and I've had a little wine!

OpenStudy (anonymous):

C=kDE/Vt

OpenStudy (anonymous):

D=(Ct/k)(V/E) (Ct/k) is constant

OpenStudy (richyw):

are you just showing someone how to do this?

OpenStudy (richyw):

or asking for help. Because you are doing it correct.

OpenStudy (anonymous):

no its my last hw problem

OpenStudy (anonymous):

got the equation dont kno how to get the answer

OpenStudy (richyw):

well if your initial equation is correct, which I didn't really check. then all your steps are correct up to here.

OpenStudy (richyw):

so change in D is \[\Delta D=D-D_0\]

OpenStudy (anonymous):

here is an example from my book

OpenStudy (richyw):

ah ok so that's a fractional change.

OpenStudy (anonymous):

shouldnt mine be 2/0.5=4 which would be 400%

OpenStudy (richyw):

yup.

OpenStudy (anonymous):

then my hw program is glitching it says it is wrong

OpenStudy (richyw):

wait did you flip your fraction around?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i just followed the example

OpenStudy (richyw):

oh right I see. yeah I have no idea what is wrong. That answer looks good to me.

OpenStudy (anonymous):

it was 300% for some reason